OpenStudy (anonymous):

Find the first 5 terms of the sequence: a1=1, an=4an-1-1

OpenStudy (anonymous):

multiply by 4, subtract 1 so for example \[a_2=4\times a_1-1=4\times 1-1=4-1=3\]

OpenStudy (anonymous):

and \[a_3=4\times a_2-1=4\times 3-1=12-1=11\]